3. Gains from trade Consider two neighboring island countries called Arcadia and Euphoria. They each have 4 million labor houIn the following table, enter each countrys production decision on the third row of the table (marked Production). SupposeCalculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relative
Answer #1

Blanks:-

1) 10/5=2

2) 16/4=4

3) Arcadia

4) Euphoria

5) 5*4=20

6) 16*4=64

7) 20-17=3

8) 64-46 = 18

13 million jeans for 39 million corn is being exchanged, so

Arcadia

Euphoria

Jeans

Rye

Jeans

Rye

(Millions of pairs)

(Millions of bushels)

(Millions of pairs)

(Millions of bushels)

Without Trade

Production

5

30

12

16

Consumption

5

30

12

16

With Trade

Production

20

0

0

64

Trade action

Export 13

Import 39

Import 13

Export 39

Consumption

7

39

13

25

Gains from Trade

Increase in Consumption

2

9

1

9